OpenStudy (anonymous):

What is the value of the expression? 15 – 10 + (200 ÷ 20)3 the 3 it a 3rd power component

OpenStudy (anonymous):

there is 200/20 ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

so than this is equal 10 what on power 3 is 1000

OpenStudy (anonymous):

what on power 3 is 1000???

OpenStudy (anonymous):

10^3 = 1000

OpenStudy (anonymous):

oh yeah ik

OpenStudy (anonymous):

then do 15-10 then add?

OpenStudy (anonymous):

it =1005

OpenStudy (anonymous):

no after 10 is divide sign ?

OpenStudy (anonymous):

??? the equation says 15 – 10 + (200 ÷ 20)^3

OpenStudy (anonymous):

next to 10 there is plus or divide ?

OpenStudy (anonymous):

+ plus

OpenStudy (anonymous):

and between 200 and 20 what is divide ?

OpenStudy (anonymous):

ok its 15-10 plus (200 divided by 20)to the 3rd power

OpenStudy (anonymous):

so than 15-10=5 and 10^3 =1000 5+1000=1005 so than you are right how you have wrote above

OpenStudy (anonymous):

thank you can yo medal me 2?

OpenStudy (anonymous):

thanks i have more questions

OpenStudy (anonymous):

ok was my pleasure good luck

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

post them on the open question column and i will try help you than you will tag me there